# Machine learning of the Ising model on a spherical Fibonacci lattice

### v2

@@ Abstract
We investigate the Ising model on a spherical surface, utilizing a Fibonacci lattice to approximate uniform coverage. This setup poses challenges in achieving consistent lattice distribution across the sphere for comparison with planar models. We employ Monte Carlo simulations and graph convolutional networks (GCNs) to study spin configurations across a range of temperatures and to determine phase transition temperatures. The Fibonacci lattice, despite its uniformity, contains irregular sites that influence spin behavior. In the ferromagnetic case, sites with fewer neighbors exhibit a higher tendency for spin flips at low temperatures, though this effect weakens as temperature increases, leading to a higher phase transition temperature than in the planar Ising model. In the antiferromagnetic case, lattice irregularities prevent the total energy from reaching its minimum at zero temperature, highlighting the role of curvature and connectivity in shaping interactions. Phase transition temperatures are derived through specific heat, magnetic susceptibility analysis and GCN predictions, yielding $T_{c}$ values for both ferromagnetic and antiferromagnetic scenarios. This work emphasizes the impact of the Fibonacci lattice’s geometric properties-namely curvature and connectivity-on spin interactions in non-planar systems, with relevance to microgravity environments.
@@ I Introduction
The Ising model is perhaps the simplest statistical spin model, yet it exhibits a wealth of physical phenomena and plays a crucial role across multiple fields of physics. For example, site disorder and tunable quantum fluctuations give rise to a rich variety of ground states[1, 2, 3, 4, 5, 6, 7, 8]. The two-dimensional Ising model is among the simplest statistical models to demonstrate a phase transition. Experimentally, the Ising model can be realized using cold-atom quantum simulators [9, 10, 11, 12, 13]. Recently, the rapid advancement of space-based technology has driven experimental efforts to confine ultracold atoms on surfaces of various shapes in microgravity [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Among these adjustable shapes, the sphere or spherical bubble trap has garnered significant research interest, with numerous studies dedicated to exploring its geometric effects on cold atomic gases. In this paper, we focus on the Ising model confined to a spherical surface. Given the fundamental nature of the Ising model, its analysis is crucial for understanding the physical characteristics of spherical atomic gases in a microgravity environment.
@@
The first challenge of this problem is to cover the spherical surface as uniformly as possible. Only in this way can we reliably compare the results with those of the planar square lattice. This requirement excludes the traditional latitude-longitude lattice, as the density of sites near the poles is significantly higher than anywhere else. The most suitable candidate is the Fibonacci lattice, which is essentially the most uniform spherical lattice. We have previously applied it to study the spherical XY model [25], yielding intriguing results about how vortex distribution is significantly influenced by spherical topology. It is important to note that the Fibonacci lattice is not perfectly uniform, which complicates the analytical study of the Ising model on the Fibonacci sphere. However, by utilizing Monte Carlo (MC) simulations and machine learning techniques, we can determine the spin configurations at various temperatures and identify the phase-transition temperature of the spherical Ising model.
@@
The second challenge arises from the irregular nature of the Fibonacci lattice, which lacks an image-like structure. Machine learning techniques have been widely used to identify phase transitions in statistical physics models, such as the Ising model. Notably, Carrasquilla and Melko [26] demonstrated the effectiveness of supervised learning with convolutional neural networks (CNNs) in classifying ordered and disordered phases in the two-dimensional square-lattice Ising model, successfully determining the critical temperature from labeled configurations. However, this approach cannot be directly applied to the Fibonacci lattice due to its non-uniform geometry. To address this, we extend the paradigm of machine learning for phase transition detection to the irregular geometry of the spherical Fibonacci lattice. Specifically, we employ graph convolutional networks (GCNs), a natural generalization of CNNs for non-Euclidean structures, which we have previously applied to study the XY model on a spherical Fibonacci lattice [25].
@@
Recently, a study focused on the ferromagnetic Ising model on a Fibonacci-triangulated sphere found that the model exhibits a critical temperature slightly lower than that of a planar triangular lattice [27]. In this paper, we will concentrate on the mostly quadrangulated Fibonacci lattice for both ferromagnetic and antiferromagnetic Ising models, where neighbor interactions are determined by a cutoff radius $r_{c}$. Through the application of specific heat analysis and GCNs, we determined the phase transition temperatures for both situations, highlighting the influence of geometric properties on spin interactions.
@@
The rest of this paper is organized as follows: In Section II, we construct the spherical Fibonacci lattice points, ensuring that the number of lattice sites with four nearest neighbors is maximized by selecting an appropriate spin-interaction cutoff radius $r_{c}$. Sections III and IV present simulations of the spherical ferromagnetic and antiferromagnetic Ising models, respectively, using MC algorithms and GCNs. These sections include graphs illustrating spin configurations, energy, entropy and specific heat at various temperatures, as well as the phase transition temperatures for both models. Finally, Section V concludes the paper.
@@ II Spherical Ising Model
The spins of a Fibonacci Ising model are located on a Fibonacci lattice on a spherical surface. The position of each lattice site in Cartesian coordinates is given by: [28, 29, 30]
@@ III.1 Spin configurations
We first consider the case with $J>0$, in which neighboring sites tend to align with the same spin orientation. This is referred to as the spherical ferromagnetic Ising model. When mapping the Ising model on a spherical Fibonacci lattice, $r_{c}$ serves as a cutoff range of the interaction. We set $J=1$, $N=1000$ and $r_{c}/R=0.1298$, and use the MC techniques to obtain samples of spin configurations. We randomly flip the spins in the system and accept or reject state transitions based on the Metropolis criterion. Specifically, if the energy of the system decreases after a flip, the flip is accepted; otherwise, it is accepted only with a certain probability, known as the acceptance probability. As the temperature decreases, the acceptance probability gradually decreases and eventually reaches equilibrium.
@@
Note that the Fibonacci lattice is inhomogeneous, with varying probabilities of spin flips at different sites, which is crucial for understanding stable spin patterns. For instance, a site with three neighbors incurs an energy cost of $6J$ to flip its spin, while a site with five neighbors incurs a cost of $10J$. Consequently, the probability of a spin-flip is relatively higher for the site with three neighbors. In the left subpanel of the top row in Fig.4, we present a view of the Fibonacci lattice from a particular direction at $T/J=2.0$. It is evident that among the sites where the spin is flipped, at least 6 out of 76 sites have three neighbors. In contrast, only 11 out of 850 sites with four neighbors exhibit spin flips. This indicates that the proportion of spin flips is relatively significant for sites with three neighbors. We also statistically calculate the average spin-flip probabilities $P_{3}$, $P_{4}$ and $P_{5}$ for sites with 3, 4 and 5 neighbours, respectively, across different stable spin configurations at various temperatures. The results are illustrated in the bottom panel of Fig.4, with black squares representing $P_{3}$, red disks representing $P_{4}$, and blue triangles representing $P_{5}$. To better illustrate the temperature-dependent changes in the relative spin-flip probabilities for different neighboring sites, we display the ranking of $P_{3}$, $P_{4}$ and $P_{5}$ as a function of temperature in the right subpanel of the top row in Fig.4, with black, red and blue squares representing $P_{3}$, $P_{4}$, and $P_{5}$ respectively. The numbers on the left, ranging from 1 to 3, indicate the relative ranking of the values rather than their specific numerical values. Specifically, we perform ten simulations at each temperature. For each simulation, we count the number of spin flips occurring at the sites with 3, 4, and 5 neighbours in a stable spin configuration. We then compute the ratio of these flips to the total number of sites for each type to obtain the flip ratio. Finally, we average the results from the ten simulations at each temperature. At low temperatures, $P_{3}>P_{4}>P_{5}$, reflecting that sites with fewer neighbors are more prone to spin flips. As the temperature increases to $T/J\gtrsim 2.34$, thermal fluctuations become more significant relative to the energy cost of spin flips, leading to a disruption in the ordering of $P_{3}$, $P_{4}$ and $P_{5}$. The inset of the bottom panel displays that $P_{4}>P_{5}>P_{3}$ at $T/J=2.34$. This effect has a significant impact on the phase transition temperature, as will be demonstrated later.
@@
Figure 5: (Top panel) Total energy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Middle panel) Pecific heat as a function of temperature, calculated using $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$. (Bottom panel) Specific heat as a function of temperature, calculated using Eq.(3).
@@ III.2 Determining $T_{c}$ via the behaviors of the specific heat and the magnetic susceptibility
Next, we aim to determine the phase transition temperature $T_{c}$ of the spherical Fibonacci Ising model. To achieve this, we employ three different approaches. The first one is straightforward but relatively approximate: We estimate $T_{c}$ by analyzing the behavior of the specific heat $C_{V}$ using three different methods for cross-validation. The second one involves estimating $T_{c}$ by analyzing the magnetization $M$ and magnetic susceptibility $\chi$. The third one offers greater accuracy and involves using graph convolutional networks (GCN) to refine our estimation of $T_{c}$.
Our first method to estimate the specific heat utilizes the formula $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$. Given a stable spin configuration, the total energy $E$ of the system is calculated via Eq.(1). We then average the total energy over 10 stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in the top panel of Fig.5. To estimate the specific heat, we use the spine functions to fit the $E$ versus $T$ curve, resulting in a function $E(T)$. We subsequently compute the derivative of this function to estimate $C_{V}$ and plot $C_{V}$ vs $T$ in the middle panel of Fig.5. As depicted in the figure, $C_{V}$ exhibits a singular behavior at $T/J\approx 2.29$, which can be considered as an estimate of $T_{c}$. Interestingly, this value is very close to the critical temperature of the two-dimensional Ising model on a square lattice, which is $T^{\Box}_{c}/J=2.269$ [31]. To quantify the uncertainty in this estimation of $T_{c}$, we apply a nonparametric Bootstrap resampling approach [32, 33], which has recently been applied to the Ising model [34]. Further details of the method are provided in Appendix A. The bootstrap iteration is performed $B=1000$ times, yielding a more precise estimate of $T_{c}$, namely $T_{c}=2.279\pm 0.008J$.
@@
where the Boltzmann constant is set to 1 in natural units. Here, the energy $E$ is obtained using the previous method, and its absolute value is quite large. Therefore, when calculating $C_{V}$ by using Eq.(3), the squared terms become even larger, which can significantly amplify the oscillations of $E$. However, the results from the previous method may be influenced by the curve fitting technique. In contrast, this method does not require calculating the slope of the fitted curve, thereby avoiding errors caused by different fitting approaches. In the bottom panel of Fig.5, we present our calculations based on Eq.(3), where each data point is averaged over 10 stable spin patterns too. As expected, the value of $C_{V}$ exhibits significant oscillations as the temperature changes. However, there is also a distinct singularity around $T\approx 2.28J$. We further attempt to fit the curve using a Gaussian function and observe that the peak is indeed located around $2.28J$. To improve the accuracy of our results, we apply the Bootstrap error analysis, which refines the estimation of $T_{c}$ to $2.335\pm 0.292J$. Although this method introduces a larger uncertainty, the confidence interval of the results still overlaps with that obtained from the previous method.
Figure 6: (Top panel) Entropy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Bottom panel) Specific heat as a function of temperature, calculated using $C_{V}=T\left(\frac{\partial S}{\partial T}\right)_{V}$. The inset shows the logarithmic fitting of data near $T_{c}$
[…]
Figure 7: (Top panel) Magnetization as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Bottom panel) Magnetic susceptibility as a function of temperature, calculated using $\chi=\frac{\langle M^{2}\rangle-\langle M\rangle^{2}}{NT}$.
@@
Through methods similar to those used for analyzing the temperature dependence of energy and specific heat, we average the total magnetization and susceptibility over 10 stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in Fig.7, where the top panel shows the variation of the magnetization with temperature, and the bottom panel shows the variation of the susceptibility with temperature. As depicted in the figure, the susceptibility $\chi$ exhibits a singular behavior at $T/J\approx 2.33$, which can be considered as an estimate of $T_{c}$. We also apply the Bootstrap error analysis to improve the accuracy of the results, which refines the estimation of $T_{c}$ to $2.316\pm 0.053J$, which is close to the estimates of $T_{c}$ obtained from the specific heat analysis.
III.3 Determining $T_{c}$ via the method of machine learning

### v3

@@ Abstract
We investigate the Ising model on a spherical surface, utilizing a Fibonacci lattice to approximate uniform coverage. This setup poses challenges in achieving consistent lattice distribution across the sphere for comparison with planar models. We employ Monte Carlo simulations, principal component analysis (PCA), graph convolutional networks (GCNs) to study spin configurations across a range of temperatures and to determine phase transition temperatures. The Fibonacci lattice, despite its uniformity, contains irregular sites that influence spin behavior. In the ferromagnetic case, sites with fewer neighbors exhibit a higher tendency for spin flips at low temperatures, though this effect weakens as temperature increases, leading to a higher phase transition temperature than in the planar Ising model. In the antiferromagnetic case, lattice irregularities induce geometric frustration, resulting in highly degenerate ground states and the phase transition temperature lower than the planar square lattice. Phase transition temperatures are derived through specific heat, magnetic susceptibility analysis and GCNs predictions, yielding $T_{c}$ values for both ferromagnetic and antiferromagnetic scenarios. This work emphasizes the impact of the Fibonacci lattice’s geometric properties-namely curvature and connectivity-on spin interactions in non-planar systems, with relevance to microgravity environments.
@@ I Introduction
The Ising model is perhaps the simplest statistical spin model, yet it exhibits a wealth of physical phenomena and plays a crucial role across multiple fields of physics. Experimentally, the Ising model can be realized using cold-atom quantum simulators Elliott et al. (1970); Orth et al. (2008); Labuhn et al. (2016); Bernien et al. (2017); Parker et al. (2013). Recently, the rapid advancement of space-based technology has driven experimental efforts to confine ultracold atoms on surfaces of various shapes in microgravity Lundblad et al. (2019a); Tononi et al. (2020a); Carollo et al. (2022); Tononi et al. (2022, 2020b); Tononi and Salasnich (2019); Lundblad et al. (2019b); Bereta et al. (2021); van Zoest et al. (2010); Becker et al. (2018). Among these adjustable shapes, the sphere or spherical bubble trap has garnered significant research interest, with numerous studies dedicated to exploring its geometric effects on cold atomic gases. In this paper, we focus on the Ising model confined to a spherical surface. Given the fundamental nature of the Ising model, its analysis is crucial for understanding the physical characteristics of spherical atomic gases.
@@
The first challenge of this problem is to cover the spherical surface as uniformly as possible. Only in this way can we reliably compare the results with those of the planar square lattice. This requirement excludes the traditional latitude-longitude lattice, as the density of sites near the poles is significantly higher than anywhere else. The most suitable candidate is the Fibonacci lattice, which is essentially the most uniform spherical lattice. It is important to note that the Fibonacci lattice is not perfectly uniform, which complicates the analytical study of the Ising model on the Fibonacci sphere. However, by utilizing Monte Carlo (MC) simulations and machine learning techniques, we can determine the spin configurations at various temperatures and identify the phase-transition temperature of the spherical Ising model.
@@
The second challenge arises from the irregular nature of the Fibonacci lattice, which lacks an image-like structure. Machine learning techniques have been widely used to identify phase transitions in statistical physics models, such as the Ising model. Notably, Carrasquilla and Melko Carrasquilla and Melko (2017) demonstrated the effectiveness of supervised learning with convolutional neural networks (CNNs) in classifying ordered and disordered phases in the two-dimensional square-lattice Ising model, successfully determining the critical temperature from labeled configurations. However, this approach cannot be directly applied to the Fibonacci lattice due to its non-uniform geometry. To address this, we extend the paradigm of machine learning for phase transition detection to the irregular geometry of the spherical Fibonacci lattice. Specifically, we employ graph convolutional networks (GCNs), a natural generalization of CNNs for non-Euclidean structures.
In our previous work, we applied the GCNs approach to study the XY model on the spherical Fibonacci lattice Song et al. (2022), yielding intriguing results about how vortex distribution is significantly influenced by spherical topology. This demonstrates that the method is well suited for handling continuous spin systems and complex non-Euclidean geometries.
@@
Recently, a study focused on the ferromagnetic Ising model on a Fibonacci-triangulated sphere found that the model exhibits a critical temperature slightly lower than that of a planar triangular lattice Pochinok et al. (2023). In this paper, we will concentrate on the mostly quadrangulated Fibonacci lattice for both ferromagnetic and antiferromagnetic Ising models, where neighbor interactions are determined by a cutoff radius $r_{c}$. Through the application of specific heat analysis and GCNs, we determined the phase transition temperatures for both situations, highlighting the influence of geometric properties on spin interactions.
@@
The rest of this paper is organized as follows: In Section II, we construct the spherical Fibonacci lattice points, ensuring that the number of lattice sites with four nearest neighbors is maximized by selecting an appropriate spin-interaction cutoff radius $r_{c}$. Sections III and IV present simulations of the spherical ferromagnetic and antiferromagnetic Ising models, respectively, using MC algorithms, PCA and GCNs. These sections include graphs illustrating spin configurations, energy, specific heat, magnetization and magnetic susceptibility at various temperatures, as well as the phase transition temperatures for both models. Finally, Section V concludes the paper.
@@ II Spherical Ising Model
The spins of a Fibonacci Ising model are located on a Fibonacci lattice on a spherical surface. The position of each lattice site in Cartesian coordinates is given by: Gonzalez (2010); Hannay and Nye (2004); Keinert et al. (2015)
@@ III.1 Spin configurations
We first consider the case with $J>0$, in which neighboring sites tend to align with the same spin orientation. This is referred to as the spherical ferromagnetic Ising model. When mapping the Ising model on a spherical Fibonacci lattice, $r_{c}$ serves as a cutoff range of the interaction. We set $J=1$, $N=1000$, and $r_{c}/R=0.1298$. and use the MC techniques to generate the spin configurations. Specifically, a spin is randomly selected and proposed to flip, and the Metropolis criterion is applied to decide whether the flip is accepted Metropolis et al. (1953); Lowry (1964); Lang et al. (1973). If the flip lowers the system energy, it is always accepted; otherwise, it is accepted with a certain acceptance probability defined as
$P_{\text{accept}}=\min\left(1,\mathrm{e}^{-\Delta E/k_{B}T}\right),$ (3)
where $\Delta E$ is the energy change caused by the flip, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. As the temperature decreases, the probability of accepting energy-increasing flips becomes smaller, reflecting the tendency of the system to remain in low-energy states and suppress disordered configurations. In the zero temperature limit, only energy-lowering flips are allowed, and the system eventually converges to the ground state. At finite temperature, after a sufficient number of spin-flip attempts, the statistical distribution of spin configurations converges to the Boltzmann distribution,
$P(\{s_{i}\})\propto e^{-E(\{s_{i}\})/k_{B}T}.$ (4)
At this stage, the acceptance probability no longer changes with the number of Monte Carlo steps but instead fluctuates around a stable mean value, indicating that the system has reached a statistical steady state, i.e., thermal equilibrium.
@@
Note that the Fibonacci lattice is inhomogeneous, with varying probabilities of spin flips at different sites, which is crucial for understanding stable spin patterns. For instance, a site with three neighbors incurs an energy cost of $6J$ to flip its spin, while a site with five neighbors incurs a cost of $10J$. Consequently, the probability of a spin-flip is relatively higher for the site with three neighbors. In the left subpanel of the top row in Fig.4, we present a view of the Fibonacci lattice from a particular direction at $T/J=2.0$. It is evident that among the sites where the spin is flipped, at least 6 out of 76 sites have three neighbors. In contrast, only 11 out of 850 sites with four neighbors exhibit spin flips. This indicates that the proportion of spin flips is relatively significant for sites with three neighbors. We also statistically calculate the average spin-flip probabilities $P_{3}$, $P_{4}$ and $P_{5}$ for sites with 3, 4 and 5 neighbours, respectively, across different stable spin configurations at various temperatures. The results are illustrated in the bottom panel of Fig.4, with black squares representing $P_{3}$, red disks representing $P_{4}$, and blue triangles representing $P_{5}$. To better illustrate the temperature-dependent changes in the relative spin-flip probabilities for different neighboring sites, we display the ranking of $P_{3}$, $P_{4}$ and $P_{5}$ as a function of temperature in the right subpanel of the top row in Fig.4, with black, red and blue squares representing $P_{3}$, $P_{4}$, and $P_{5}$ respectively. The numbers on the left, ranging from 1 to 3, indicate the relative ranking of the values rather than their specific numerical values.
Specifically, we perform 100 independent Monte Carlo simulations at each temperature to obtain 100 independent spin configurations. Each simulation is initialized with a random spin configuration and evolved according to the Metropolis criterion for $5\times 10^{5}$ steps to ensure thermal equilibration. In each simulation, we count the number of spin flips occurring at sites with 3, 4, and 5 neighbours in the stable spin configurations and divide these counts by the total number of sites of the corresponding type to obtain the flip ratios. Finally, the results from the 100 simulations at each temperature are averaged to obtain statistically reliable spin-flip probabilities. At low temperatures, $P_{3}>P_{4}>P_{5}$, reflecting that sites with fewer neighbors are more prone to spin flips. As the temperature increases to $T/J\gtrsim 2.34$, thermal fluctuations become more significant relative to the energy cost of spin flips, leading to a disruption in the ordering of $P_{3}$, $P_{4}$ and $P_{5}$. The bottom panel displays that $P_{4}>P_{5}>P_{3}$ at $T/J=2.34$. This effect has a significant impact on the phase transition temperature, as will be demonstrated later.
@@
Figure 5: (Top panel) Total energy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Middle panel) Specific heat as a function of temperature, calculated using $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$. (Bottom panel) Specific heat as a function of temperature, calculated using Eq.(5). The black dashed lines represent the raw Monte Carlo data points, and the red solid line shows the Gaussian fit around the peak.
@@ III.2 Determining $T_{c}$ via the behaviors of the specific heat and the magnetic susceptibility
Next, we aim to determine the phase transition temperature $T_{c}$ of the spherical Fibonacci Ising model. To achieve this, we employ three different approaches. The first one is straightforward but relatively approximate: We estimate $T_{c}$ by analyzing the behavior of the specific heat $C_{V}$ using three different methods for cross-validation. The second one involves estimating $T_{c}$ by analyzing the magnetization $M$ and magnetic susceptibility $\chi$. The third one offers greater accuracy and involves using graph convolutional networks (GCNs) to refine our estimation of $T_{c}$.
Our first method to estimate the specific heat utilizes the formula $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$. Given a stable spin configuration, the total energy $E$ of the system is calculated via Eq.(1). We then average the total energy over 100 stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in the top panel of Fig.5. To estimate the specific heat, we use the spline functions to fit the $E$ versus $T$ curve, resulting in a function $E(T)$. We subsequently compute the derivative of this function to estimate $C_{V}$ and plot $C_{V}$ vs $T$ in the middle panel of Fig.5. As depicted in the figure, $C_{V}$ exhibits a peak behavior at $T/J\approx 2.28$, which can be considered as an estimate of $T_{c}$. Interestingly, this value is very close to the critical temperature of the two dimensional Ising model on a square lattice, which is $T^{\Box}_{c}/J=2.269$ Kramers and Wannier (1941); Onsager (1944); Kaufman (1949). To quantify the uncertainty in this estimation of $T_{c}$, we apply a nonparametric Bootstrap resampling approach Efron (1992); Kawano and Higuchi (1995), which has recently been applied to the Ising model Lee (2022). Further details of the method are provided in Appendix A. The bootstrap iteration is performed $B=1000$ times, yielding a more precise estimate of $T_{c}$, namely $T_{c}/J=2.279\pm 0.005$.
@@
where the Boltzmann constant is set to 1 in natural units. Here, the energy $E$ is obtained using the previous method, and its absolute value is quite large. Therefore, when calculating $C_{V}$ by using Eq.(5), the squared terms become even larger, which can significantly amplify the oscillations of $E$. However, the results from the previous method may be influenced by the curve fitting technique. In contrast, this method does not require calculating the slope of the fitted curve, thereby avoiding errors caused by different fitting approaches. In the bottom panel of Fig.5, we present our calculations based on Eq.(5), where each data point is averaged over 100 stable spin patterns too. As expected, the value of $C_{V}$ exhibits significant oscillations as the temperature changes. However, there is also a distinct maximum around $T/J\approx 2.313$. We further attempt to fit the curve using a Gaussian function and observe that the peak is indeed located around $2.313J$. To improve the accuracy of our results, we apply the Bootstrap error analysis, which refines the estimation of $T_{c}/J$ to $2.31\pm 0.007$. Although this method introduces a larger uncertainty, the confidence interval of the results still overlaps with that obtained from the previous method.
Figure 6: Top panel) Magnetization as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Bottom panel) Magnetic susceptibility as a function of temperature, calculated using $\chi=\frac{\langle M^{2}\rangle-\langle M\rangle^{2}}{NT}$. The black dashed lines represent the raw Monte Carlo data points, and the red solid line shows the Gaussian fit around the peak.
@@
Through methods similar to those used for analyzing the temperature dependence of energy and specific heat, we average the total magnetization and susceptibility over 100 stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in Fig.6, where the top panel shows the variation of the magnetization with temperature, and the bottom panel shows the variation of the susceptibility with temperature. As depicted in the figure, the susceptibility $\chi$ exhibits a peak behavior at $T/J\approx 2.369$, which can be considered as an estimate of $T_{c}$. We also apply the Bootstrap error analysis to improve the accuracy of the results, which refines the estimation of $T_{c}/J$ to $2.370\pm 0.005$, which is close to the estimates of $T_{c}$ obtained from the specific heat analysis.
III.3 Determining $T_{c}$ via the method of GCNs